OpenStudy (anonymous):

I need help with complex numbers, can someone please guide me through this prob?

OpenStudy (anonymous):

Find the value of k for the quotient \[\frac{ 2 - ki }{ k - i }\] if it is: 1: A pure imaginary number. 2: A real number.

OpenStudy (unklerhaukus):

if \(k\) is purely imaginary, then for a real constant \(b\); \(k=ib\)

OpenStudy (unklerhaukus):

\[\frac{2-ki}{k-i}\implies\frac{2-(ib)i}{(bi)-i}\] simplify this

OpenStudy (anonymous):

Oh okay, so we're assuming this divides evenly with no remainder?

OpenStudy (unklerhaukus):

the quotient is the whole number if times it divides, the remainder is forgotten

OpenStudy (anonymous):

i c

OpenStudy (anonymous):

Correct and I see that "pure imaginary number" means a complex number with no "real" part. So a complex number is a + bi where a is the real part and bi is imaginary term So from that you got k = bi

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

so for starters ... 2 - bi^2 right?

OpenStudy (unklerhaukus):

for the numerator: yeah, and use i^2 = -1

OpenStudy (anonymous):

\[\frac{ 2 - bi ^{2} }{ i(b-1) }\]

OpenStudy (anonymous):

\[\frac{ 2+b }{ i (b-1) }\]

OpenStudy (anonymous):

we are supposed to find a value for k, so we need to figure out what b in bi is let me think about it ...

OpenStudy (unklerhaukus):

what do you get if you multiply this by \(\frac ii\)

OpenStudy (anonymous):

Hmm something is still missing from the problem ... I don't think you can solve if you don't know what the fractions equals It's not an equation and k is supposed to be a number so you can't do long division k is a number, it's supposed to divide to be another number Like if you have 24/4 = 6 You have to get the fraction equal to a number Right now we don't know anything about what the fraction equals so you can't solve

OpenStudy (anonymous):

well, I don't think you need the fraction to equal to something in order for it to solve :/

OpenStudy (anonymous):

whoever wrote the question just stated it weirdly I reckon

OpenStudy (anonymous):

I notice we started from a quotient but it has imaginary terms in it I think they want me to change that into a + bi form first So to do that, we can multiply by the conjugate of the bottom ... \[\frac{ (2 - ki) }{ k-i }\]

OpenStudy (anonymous):

I'll try to figure something out Thanks for your help, @UnkleRhaukus :)

OpenStudy (unklerhaukus):

(something about the question isn't making complete sense to me either)

OpenStudy (anonymous):

ok so like for this whole time I was talking to myself trying to get this prob to somehow interpret but aye what can u do with a dyslexic person like me merp, ty anyways this prob has got sum issues srsly imma go consult a teacher and he/she shall deal with it cause am going nuts about it ok bye

OpenStudy (unklerhaukus):

okbye